🚀 go-pugleaf

RetroBBS NetNews Server

Inspired by RockSolid Light RIP Retro Guy

Thread View: sci.astro.hubble
1 messages
1 total messages Started by hst-status Tue, 22 Dec 1992 15:15
HST Status Report, 12/17 - 12/18
#14
Author: hst-status
Date: Tue, 22 Dec 1992 15:15
98 lines
3911 bytes
                            HUBBLE SPACE TELESCOPE
                               DAILY REPORT #787

PERIOD COVERED:  0800 EST 12/17/92 - 0800 EST 12/18/92

Daily Status Report as of 353/1200Z.

1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED:

   1.1 Complete One Set of GHRS/2 3482 (Nature of In-Falling Matter on Star
       Beta Pictoris) and Initiate a Second.

       The High Resolution Spectrograph was used to take a series of 19
       ultraviolet spectral observations of the gaseous disk around Beta
       Pictoris (aka, HD39060) looking for cometary like objects which
       might be falling into the disk and being evaporated. The onboard
       target acquisition was successful. The wave calibrations looked
       okay, and all the spectral observations were reported to look good.
       A second set of observations are currently under way, and no
       problems have been reported. 
 
   1.2 Complete FOS/RD 3840 (Abundance of Carbon, Nitrogen, and Oxygen in 
       ST)

       The Faint Object Spectrograph was used to make four ultraviolet
       spectral observations of the extra-galactic H II region IZW18. The
       binary search found the target in group eight, and looked okay.
       All of the spectral observations were reported to be okay. The
       second confirmation image also looked okay. 

2.0 FLIGHT OPERATIONS SUMMARY:

    2.1 Guide Star Acquisitions: 

                            Scheduled Acquisitions : 2
                                        Successful : 2
                          Scheduled Reacquisitions : 9
                                        Successful : 9
                                    Losses of Lock : 0

   2.2 FHST Updates: 

                              Scheduled : 7
                             Successful : 7

   2.3 Significant HSTARs: None

   2.4 Operations Note: 

       The First Servicing Mission Internal Simulation #2 was successfully
       completed yesterday, Thursday, Dec. 17. This simulation was intended
       to provided the HST Servicing Mission team with functional training
       to ensure operational readiness. This simulation went well and no
       major problems were reported. 

3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD:

   3.1 Complete Six Sets of GHRS/2 3482 (Nature and In-Falling of Matter
       on Star Beta Pictoris) 

   3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871

   3.3 Complete PC 4093 (Survey of Nuclear Regions of Markarian Galaxies)

   3.4 Complete WFC/PC 4109 (WF/PC Calibration:Rapid Internal Monitor)

   3.5 Complete Two Sets of GHRS/2 4068 (GHRS Detector Calibrations)

   3.6 Complete PC 4140 (WF/PC Calibration: Rapid Internal Monitor Part 2)

   3.7 Complete PC 4163 (Evolution of Solar Nebulae Surrounding Pre-Main 
       Sequence)

   3.8 Complete GHRS/2 4065 (Dark Noise Monitoring HRS)

   3.9 Complete PC 3551 (Elliptical Galaxies with Kinematically Distinct 
       Nuclei)

   3.10 Complete FOC/96 4096 (Far Ultraviolet Spectra of High redshift 
       Quasars)

   3.11 Complete FOS/BL 3816 (White Dwarf Stars)

   3.12 Complete WFC 4105 (HST Medium Deep Survey,High Latitude)

   3.13 Complete WFC 4713 (Internal Delta Flats)

   3.14 Complete GHRS/2 2851 (Pulse Height Analysis (PHA))

4.0 ANOMALY STATUS:
	
   4.1 Significant New Anomalies: None

   4.2 Significant Unresolved Anomalies: None
 
 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time
 
                          -END-
Thread Navigation

This is a paginated view of messages in the thread with full content displayed inline.

Messages are displayed in chronological order, with the original post highlighted in green.

Use pagination controls to navigate through all messages in large threads.

Back to All Threads